# Quasicrystals as an intermediate form of matter between crystalline and amorphous solids

### v1

@@ Abstract
Quasicrystals have been observed in a variety of materials ranging from metal alloys to block copolymers and represent an “intermediate” form of matter between crystals and amorphous materials (glasses and liquids) in that their structural and dynamical properties can not readily described in terms of conventional solid-state models of liquids and solids. In the present work, we present a comprehensive analysis of basic thermodynamic and dynamic properties of quasicrystals to better understand the nature of the atomic motion underlying diffusion and structural relaxation in these materials. As our model system, we investigate a dodecagonal quasicrystal using molecular dynamics (MD) simulations in two dimensions (2D), subject to periodic boundary conditions. We observe a two-stage relaxation dynamics in the self-intermediate scattering function $F_{s}(k,t)$ of our quasicrystal material involving a fast $\beta$-relaxation on a ps timescale and $\alpha$ relaxation process having a highly temperature dependent relaxation time whose activation energy varies in concert with the extent of string-like collective motion. Multi-step relaxation of the intermediate scattering function and string-like collective atomic motion have similarly been observed ubiquitously in glass-forming liquids at low temperatures and in crystalline materials at elevated temperatures where structural relaxation and diffusion are both non-Arrhenius. After examining the dynamics of our quasi-crystalline material in great detail, we conclude that its dynamics more closely resemble observations on metallic glass-forming liquids, in qualitative accord with previous neutron scattering studies.
@@ I Introduction
According to the traditional paradigm of solid-state physics, solids are divided into two types: crystals and amorphous (or non-crystalline) materials. Crystals display long-range order and they are defined by an ordered lattice with translational periodicity. They are compatible only with discrete $1$-, $2$-, $3$-, $4$-, and $6$-fold rotational symmetries and they present sharp Bragg peaks in x-ray diffraction. On the contrary, amorphous solids lack all of these properties, and in particular, they do not exhibit long-range order. In April 1982, Shechtman [1] (see also Ref. 2) discovered a crystal exhibiting $5$-fold symmetry – a type of ordered structure previously considered to be “impossible”. Despite lacking periodicity, the icosahedral phase with $5$-fold symmetry was shown to exhibit well-defined and bright Bragg peaks as for crystals, and differently from amorphous solids. The crystal that Shechtman discovered was the first example of an aperiodic crystal, or as latter termed by Levine and Steinhardt [3, 4] a “quasicrystal”, short for quasi-periodic crystal [5, 6, 7, 8]. It was recognized early by Ishimasa et al. [9] that quasicrystals represent an intermediate ordered state between crystalline and amorphous matter (see Fig. 1), which is not periodic, i.e., quasi-periodic.
Figure 1: Quasicrystals as an intermediate form of matter.
@@
Angell [10] has argued that many forms of other matter likewise qualify as being called intermediate forms of matter, e.g., proteins, plastic crystals, incommensurate structures, twisted 2D heterostructures and glass-formers undergoing fragile-to-strong glass formation, and he observed that such materials often seemed to share common thermodynamic and dynamic characteristics, even if the fundamental origin of these properties is not currently understood theoretically. In the case of quasi-crystals, neutron scattering studies have repeatedly indicated that the properties of quasi-crystalline materials and incommensurate structures often resemble those of metallic glasses rather than crystalline materials. This was first noticed for the static structure factor [11], and then recognized also for dynamical properties, connected to anomalies in the vibrational density of states (vDOS) and the low-temperature heat capacity [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Given the relatively ordered nature of quasi-crystalline materials in relation to metallic glass materials, it is then interesting to take a closer look at the dynamics of these materials and to compare with established trends for metallic glasses and heated crystalline materials to determine if this similarity to glass-forming materials also holds in relation to structural relaxation and atomic diffusion. We next review some essential properties of quasi-crystalline materials.
Mathematically, every aperiodic structure in $d$ dimensions can be obtained using projection/cut operators from a periodic structure in a higher-dimensional space (the “superspace”), which is equivalent to saying that a basis of $d$ independent unit vectors $\hat{x}_{n}$ is incomplete to describe the reciprocal lattice in a quasicrystal. The number of additional fundamental vectors needed to determine the position of the lattice points in an aperiodic structure coincides exactly with that of the extra-dimensions in the superspace [25, 26]. This additional extra-dimensional freedom has a theoreticallly predicted consequence. In quasicrystals, one expects low-energy excitations additional to phonons, termed “phasons” [27]. These predicted excitations are associated with infinitesimal rigid translations along these extra directions [28, 29, 30, 31, 32, 33, 34, 35]. But their stability and observability are outstanding questions. From hydrodynamics and effective field theory, one can formally derive that at long-wavelengths phasons are not propagating waves but rather overdamped diffusive modes [36, 37, 38, 39, 34, 35]. This situation changes for large wave-vectors, where the propagating character is restored following a telegrapher-type dynamics [40]. Both these regimes have been claimed to be observed [41, 42, 43], especially in incommensurate structures [44, 45]. The possible diffusive nature of the phason is related to the observation that rigid translations in the extra-dimensional superspace do not correspond to vibrational displacements in real space, but rather to atomic re-arrangements and particle jumps. Despite these widespread concepts, some authors [46] (see also Ref. 47 for a discussion) have questioned the microscopic origin of the phason modes and have indicated that there is no physical evidence for the existence of long wavelength phason fluctuations, despite the observation of this collective atomic motion.
On the other hand, the problem of bulk self-diffusion in quasicrystals has attracted a lot of attention in the past as a possible smoking gun for the existence of phasons. In 1993, Kalugin and Katz [48] proposed a novel mechanism for diffusion in quasicrystals based on phason dynamics. More concretely, they argued that in quasicrystals the presence of vacancies is not necessary for atomic jumps since the latter are naturally provided by the phason degrees of freedom. The temperature ($T$) dependence of the diffusivity in quasicrystals should then exhibit a deviation from a straight line in the Arrhenius plot, which is the case for vacancy-mediated diffusion in ordered crystals. A first realization of this proposal was observed by Joseph et al. [49] using Monte Carlo simulations. A two-phase diffusive dynamics, interpolating between different thermally activated regimes, was experimentally observed in Ref. 50 using radiotracer techniques (see also Ref. 51), where it was proposed that high-temperature diffusion in quasicrystals is dominated by vacancies while the low-temperature one by phason modes. The same observations were reported using NMR (Nuclear Magnetic Resonance Spectroscopy) experiments [52, 53] and quasi-elastic neutron scattering data [54] (see also Ref. 55). Phason defects have been also identified using scanning tunneling microscopy and their dynamics is related to the static stress field of dislocations [56]. Despite all these observations, it was pointed out by Dolinsek et al. [57] (see also Comment by Coddens [58]) that, although diffusive atomic jumps in quasicrystals exhibit properties compatible with the elementary excitations in the form of phason jumps, very similar atomic motion is observed also in “vacancy-ordered” bcc (body center cubic) crystals. This poses serious doubts on whether this motion is specific to quasicrystals or just a generic feature of systems that contain structural vacancies, and whether phason jumps are qualitatively different, aside from their low activation energy, in comparison to single vacancy diffusion. Additional evidence against phason-mediated diffusion has also been reported [59].
@@
From the foregoing discussion, we may expect the dynamics of quasicrystals to be distinct in some ways from the dynamics of crystalline solids, and some similarities to glass-forming liquids can be anticipated. Molecular dynamics (MD) simulations provide a suitable tool for elucidating basic thermodynamic and dynamic properties of quasicrystals. In this regard, Engel and Trebin [60] introduced the Lennard-Jones-Gauss (LJG) system which constitutes a simple model for studying quasicrystals. The potential has proven to be able to reproduce several entropically stabilized quasicrystal phases, e.g., see Ref. 61. From an atomistic level, MD simulations also provided early evidence for short-time stochastic particle dynamics in the form of phason-flips [62]. Interestingly, this can happen either with single-particle jumps but also, and more interestingly, with correlated ring-like multi-particle motion [62] or carousel-like dynamics [63]. This string-like collective dynamics is evidently another commonality between quasicrystals and glasses.
Despite the strong ongoing interest in quasicrystals from different directions, a complete description and understanding of their vibrational, thermodynamic and collective properties is absent. Moreover, a final answer to the question of whether phasons exist and which is their microscopic origin have not been reached yet. Finally, the understanding of quasicrystal dynamics from an atomistic perspective and the precise characterization of their degree of similarity with amorphous solids and liquids are still largely unexplored. The present work aims to provide a first step towards this direction by carrying out an exhaustive analysis of a 2D dodecagonal quasicrystal using MD simulations.
In the present study, we delve into a comprehensive examination of the thermodynamic properties, dynamic characteristics, and the structural origins of dynamics in quasicrystals. Our analysis is based on atomistic simulations, focusing on a model system comprising identical particles interacting in two dimensions through the LJG-type pair potential. Sec. II is dedicated to introducing the LJG-type potential, outlining the phase diagram, describing the dodecagonal quasicrystal system, and providing essential details regarding our computer simulations. In this section, we consider an exploration of the thermodynamics of a dodecagonal quasicrystal. In Sec. III, we delve into the study of relaxation mechanisms, dynamic heterogeneity, bulk self-diffusion, and the intricate interplay among these properties. At lower temperatures, we observe phason flips as part of the dynamic behavior in quasicrystals. Meanwhile, at higher temperatures, our analysis reveals temperature-dependent string-like motions within the quasicrystal. Additionally, we uncover that both individual phason flips and collective string-like motions originate from the transfer of defects within the quasicrystal. Furthermore, we introduce the concept of “string lifetime” as a important characteristic timescale in the dynamics of quasicrystals. Finally, in Sec. IV, we conclude with a comprehensive discussion of our findings. Our research offers valuable insights into the fundamental physical and mechanistic aspects governing the thermodynamic and dynamic properties of dodecagonal quasicrystals and other quasicrystalline structures.
@@ II.1 Lennard-Jones-Gauss potential
We conducted classical molecular dynamics (MD) simulations using the LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulation) software [64] to investigate the thermodynamic and dynamic properties of a 2D dodecagonal quasicrystal. Our system comprises identical particles interacting via a novel LJG-type potential, which combines the traditional Lennard-Jones potential with a secondary Gaussian well at longer distances. The LJG potential is defined as follows:
$V\left(r\right)=4{\varepsilon_{0}}\left[{{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-12}}-{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-6}}}\right]+\frac{h}{{{\sigma{\rm{G}}}\sqrt{2\pi}}}\exp\left[{-\frac{{{{\left({r-{r_{\rm{G}}}}\right)}^{2}}}}{{2\sigma_{\rm{G}}^{2}}}}\right].$ (1)
In this equation, $\varepsilon_{0}$ represents the energy unit, and $\sigma_{0}$ is the length unit. The parameter $h$, in conjunction with the standard deviation $\sigma_{\text{G}}$, dictates the peak height of the Gaussian component, while $r_{\mathrm{G}}$ denotes the peak position of the Gaussian well. After extensive experimentation, the chosen parameter values that resulted in the formation of a dodecagonal quasicrystal are $\varepsilon_{\mathrm{0}}=0.5$, $\sigma_{\mathrm{0}}=0.88$, $h=-0.874$, $r_{\mathrm{G}}=1.879$, and $\sigma_{\textrm{G}}^{2}=0.02$. All quantities are expressed in dimensionless Lennard-Jones units. Notably, with these carefully selected parameters, the LJG-type potential exhibits a double-well nature, as illustrated in Fig. 2(a). For practical purposes, we applied a cutoff to the LJG-type potential at $r=2.5$.
@@ II.2 Phase diagram
A system comprising $N=1024$ particles arranged in a square lattice is subjected to a controlled heating process, transitioning into the liquid phase at a temperature of $T=1.95$. The simulations are conducted with periodic boundary conditions applied in both spatial dimensions. Following an initial period of 50,000 molecular dynamics (MD) steps to ensure equilibration, the system undergoes a controlled linear temperature decrease to reach $T=0.001$ over a course of $3.9\times 10^{7}$ MD steps. To maintain precise control over temperature and pressure within the isothermal-isobaric ensemble, we employ two crucial techniques. The Nosé-Hoover thermostat, as introduced by Nosé [65] and further refined by Hoover [66], regulates temperature fluctuations throughout the simulations. Meanwhile, the Parrinello-Rahman barostat, developed by Parrinello and Rahman [67], ensures stable pressure conditions.
During the cooling process, the system exhibits two first-order phase transitions. The transition from phase 1 (liquid) to phase 2 (hexatic) happens at about $T_{LH}=1.78$. In order to quantitatively identify the location of the transition between these phases, and qualify the nature of phase 2, we calculate the sixfold bond-orientation order parameter [68, 69] according to
@@
where $N_{j}$ is the number of nearest neighbors of particle $j$, $\mathbf{r}_{j}$ is the position of particle $j$, and $\theta_{m}^{j}$ is the angle between the vector ($\mathbf{r}_{m}-\mathbf{r}_{j}$) and the $x$ axis. The cutoff distance for determining these nearest neighbors is set to be $1.2$, which corresponds to the bond length derived from the radial distribution function (RDF), $g(r)$, as discussed below. The average order parameter is then obtained via $\overline{\Psi}_{6}=\sum_{j=1}^{N}\psi_{6}^{j}$. The temperature dependence of $\hat{\psi}_{6}$ is shown in Fig. 3(a), which clearly indicates two thermodynamic transitions at high and intermediate temperatures.
@@
where $L$ is the length of the simulation box, and $N$ the number of atoms. The quasi-long-range bond-orientation order arises from the algebraic decay of $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-\eta_{6}}$. It is noteworthy that at the boundary between the liquid phase and the hexatic phase, $\eta_{6}$ assumes a value of 1/4. In the liquid phase, the system exhibits only short-range bond-orientation order. Consequently, the bond-orientation correlation function is more aptly described by the Ornstein-Zernike (OZ) function, characterized by $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-1/2}\exp\left(-r/\xi_{6}\right)$ [70, 71, 68]. As evident in Fig. 3(b), $g_{6}(r)/g(r)$ exhibits a gradual decay with increasing temperature. We observe a power-law decay with an exponent of $\eta_{6}=-1/4$ for $g_{6}(r)/g(r)$ approximately at $T_{LH}=1.78$, signifying the location of the liquid-hexatic transition.
The transition from phase 2 (hexatic) to phase 3 (quasicrystal) takes place at approximately $T_{HQ}=0.87$. Through a detailed comparison of the local configurations in phase 3 with experimental observations obtained via high-resolution transmission electron microscopy (TEM) [72], as illustrated in Fig. 4, we confirm that phase 3 corresponds to a monolayer dodecagonal (12-fold) symmetry quasicrystal, pointing to the relevance of our simulations to real-world quasicrystal material. Our model quasicrystal exhibits dodecagonal units and a point defect structure within a hexagonal ring configuration, both of which align with the observations made through TEM [72] and other experimental observation [26]. The agreement between simulation results and experimental observations confirms the fidelity of our model in capturing the intricate atomic arrangement of this remarkable quasicrystal phase. Further exploration of this configuration, both in direct and reciprocal space, will be the focus of our analysis in Section III.
We note that there have been many previous simulations of two dimensional crystalline materials in which corresponding transitions from a low temperature hexagonal crystalline state to an orientationally correlated hexatic state occurs upon heating, followed by an additional transition between the hexatic and liquid states at an even higher characteristic temperature. Although the well-known theories of these transitions by Kosterloitz and Thouless [73] and Halperin and Nelson [74, 75], many studies have established that these transitions can be either of second or first order in nature [76], depending on the particular form of the interparticle potential, and the presence of appreciable finite size effects can often make the assignment of transition order computationally challenging and sometimes controversial. Resolving the difference between weakly first order transitions and second order transitions can be inherently difficult [77].
There have also been studies of the collective dynamics of two dimensional systemes exhibiting crystal melting and formation of an intermediate hexatic phase [78, 79, 80]. In the present exploratory study of a model quasi-crystalline material we simply note that the thermodynamic transitions that we observe in our quasi-crystal material simulations are strikingly reminiscent to the hexagonal to hexatic and hexatic to liquid transitions established to exist in many two-dimensional crystal forming materials. Moreover, these transitions in our quasi-crystalline forming material are relatively sharp, suggestive of a weakly first order phase transitions. Given the limited size of our simulated materials, however, we do not insist on this designation of transition order and simply note that apparently well-defined thermodynamic transitions occur near certain characteristic temperatures at which the degree of material order and dynamics clearly occur. We are not aware of a previous work exploring these transitions in quasi-crystalline materials.
Previous work on crystal melting have also found a prevalence for string-like collective motion [78, 79], a phenomenon that appears superficially very similar to the observations that we report below in our simulated model quasi-crystal forming material. Despite the superficial similarities in the thermodynamic and dynamic properties of two dimensional crystal and quasi-crystal forming materials, we stop short of equating their properties. Nonetheless, the evidence of evident long-range positional correlations in our quasicrystalline material point to a greater ”similarity” to crystalline materials than glass-forming materials. However, given the somewhat contradictory indications from simulation and measurements of a ”similarity” of quasi-crystalline materials to both crystalline and glass-forming materials, we take a fresh look at which class of materials more closely resembles quasi-crystalline materials.
@@
Figure 4: The atomic structure of the monolayer dodecagonal (12-fold) quasicrystal, following relaxation at $T=0.3$, finds support from TEM observations [72].
II.3 Quasicrystal structure and dynamics
On one hand, the radial distribution function (RDF) $g(r)$ provides insights into the probability of locating a particle at a distance $r$ from the reference particle’s position in direct space. In the case of crystalline materials, $g(r)$ exhibits distinct peaks corresponding to the periodic arrangements of particles, signifying the presence of long-range order. For isotropic systems, $g(r)$ relies solely on the magnitude of the relative atomic distances and can be calculated as follows:
$g\left(r\right)=\frac{1}{2\pi\,r\,dr}\frac{1}{\rho}\left\langle\sum_{i\neq i^{\prime}}\delta\left(r-\lvert\mathbf{r}_{i}-\mathbf{r}_{i^{\prime}}\rvert\right)\right\rangle,$ (4)
where $\rho$ represents the number density.

### v2

# Atomistic mechanisms of dynamics in a two-dimensional dodecagonal quasicrystal

@@ Abstract
Quasicrystals have been observed in a variety of materials ranging from metal alloys to block copolymers. However, their structural and dynamical properties cannot be readily described in terms of conventional solid-state models of liquids and solids. We may expect the dynamics of this specific class of quasicrystalline materials to be more like glass-forming liquids in the sense of exhibiting large fluctuations in the local mobility (”dynamic heterogeneity”) and non-Arrhenius temperature dependence of relaxation and diffusion. In this work, we investigate a model dodecagonal quasicrystal material in two dimensions (2D) using molecular dynamics (MD) simulations, with a focus on heterogeneous dynamics and non-Arrhenius relaxation and diffusion. As observed in glass-forming liquids and heated crystals, we observe a two-stage relaxation dynamics in the self-intermediate scattering function $F_{s}(k,t)$ of our quasicrystal material. It involves a fast $\beta$-relaxation and $\alpha$ relaxation process having a highly temperature dependent relaxation time whose activation energy varies in concert with the extent of string-like collective motion, a phenomenon recognized to occur in glass-forming liquids at low temperatures and crystalline materials at elevated temperatures. After examining the dynamics of our dodecagonal quasicrystalline material in great detail, we conclude that the dynamics of these materials more closely resembles observations on metallic glass-forming liquids than crystalline materials.
@@ I Introduction
According to the traditional paradigm of solid-state physics, solids are divided into two types: crystals and amorphous (or non-crystalline) materials. Crystals display long-range order and they are defined by an ordered lattice with translational periodicity. They are compatible only with discrete $1$-, $2$-, $3$-, $4$-, and $6$-fold rotational symmetries and they present sharp Bragg peaks in x-ray diffraction. On the contrary, amorphous solids lack all of these properties, and, in particular, they do not exhibit long-range order. In 1982, Shechtman [1] (see also [2]) discovered a crystal exhibiting $5$-fold symmetry – a type of ordered structure previously thought to be “impossible”. Despite lacking periodicity, an icosahedral phase with $5$-fold symmetry was shown to exist and to exhibit well-defined and intense Bragg peaks as found for crystals, and distinct from amorphous solids where only broad peaks are observed associated with local ordering in the material. The crystal that Shechtman discovered was the first example of an aperiodic crystal, or as latter termed by Levine and Steinhardt [3, 4] a “quasicrystal”, short for quasi-periodic crystal [5, 6, 7, 8]. It was recognized early by Ishimasa et al. [9] that quasicrystals represent an intermediate ordered state between crystalline and amorphous matter which is not periodic.
@@
Angell [10] has argued that many forms of matter qualify as being characterized as “intermediate” forms of matter, e.g., globular proteins, plastic crystals, incommensurate structures, twisted 2D heterostructures, and glass-formers undergoing fragile-to-strong glass formation, and he observed that such materials often seemed to share common thermodynamic and dynamic characteristics, even if the fundamental origin of these properties is not currently understood theoretically. In the case of quasicrystals, neutron scattering studies have repeatedly indicated that the properties of quasicrystalline materials and incommensurate structures often resemble those of metallic glasses rather than crystalline materials. This was first noticed for the static structure factor of quasicrystalline materials [11], and then recognized also for dynamical properties, connected to anomalies in the vibrational density of states (vDOS) and the low-temperature heat capacity [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. We point out that while quasicrystals exhibit long-range orientational order as found in crystalline materials also exhibit a large configurational entropy, as in glass-forming liquids above their glass transition because of the many realizations of the quasicrystalline configurations [27]. Indeed, the remarkable thermodynamic stability of the quasicrystal materials has recently been argued to be a direct consequence of this large configurational entropy [27]. In history, the concept of entropic stabilization of quasicrystals, particularly the random tiling hypothesis, has been explored for nearly 30 years [28, 29].
Given the relatively ordered nature of quasicrystalline materials in relation to metallic glass materials, it is then interesting to take a closer look at the dynamics of these materials and to compare with established trends for metallic glasses and heated crystalline materials to determine if this similarity to glass-forming materials also holds in relation to structural relaxation and atomic diffusion.
Since the discovery of quasicrystals there has been avalanche of scientific activity aimed at discovering real materials of this kind and in theoretically understanding the thermodynamics and dynamic properties of quasicrystalline materials which are now known to occur in many forms. Engel [30] made a pioneering study of the thermodynamics and dynamics of decagonal and dodecagonal quasicrystals, the two most prevalent types of quasicrystals that occur in near two-dimensional (2D) materials. The dynamic behavior of quasicrystals is typically described by two elementary excitations: phonon and phason. Phonon describes the atomic oscillatory motion around equilibrium positions. Phason modes correspond to the local rearrangements of atoms. The phason modes and the associated dynamics involving flips have been studied extensively in the literature [31, 32]. At the continuum level, phason modes are diffusive in nature. The elementary process of a phason mode is ”phason flip” on the atomistic scale. A phason flip entails the overcoming of an energy barrier, resulting in the transformation of a local configuration into another with a comparable topological structure and proximate energy level [30]. The decagonal type of quasicrystal was found to exhibit a relative local particle displacement dynamics occurring through phason flips, which led to a nearly Arrhenius temperature dependence of diffusion in this type of liquid crystal. However, collective particle exchange motion and non-Arrhenius diffusion were found to be more prevalent in dodecagonal quasicrystals, a class of quasicrystalline materials that is more prevalent in soft matter materials such as polymers than metallic materials. [33, 34]. Damasceno [35] have reviewed numerous observations on these model quasicrystalline materials following their initial study.
This study employs a comprehensive approach to investigate the thermodynamic properties and dynamic behavior of 2D dodecagonal quasicrystal, while establishing a systematic comparative framework to identify structural and dynamical parallels with glass-forming liquids and heated crystals. The experimental claims that quasicrystals resemble glass-forming liquids above really do not answer these questions. A careful analysis of the heterogeneous and collective dynamics of the quasicrystal is required, since as we shall see below, heated crystals and glass-forming liquids have many common features in their dynamics that make answering this question more difficult than one might expect.
@@
Molecular dynamics (MD) simulations provide a suitable tool for elucidating basic thermodynamic and dynamic properties of quasicrystals. In this regard, Engel and Trebin [36] introduced the Lennard-Jones-Gauss (LJG) as a model for 2D quasicrystals. The potential has proven to be able to reproduce several entropically stabilized quasicrystal phases, e.g., see Ref. [37]. From an atomistic level, MD simulations also provided early evidence for short-time stochastic particle dynamics in the form of phason-flips [30]. Interestingly, this can happen either with single-particle jumps, but also correlated multi-particle exchange events [30, 38]. This string-like collective dynamics is evidently suggestive of a link between quasicrystals and glasses, where this type of particle exchange motion is ubiquitous.
Since there has been extensive previous work on the LJG model for molecular parameters and temperature values leading to dodecagonal quasicrystalline materials, we briefly summarize thermodynamic properties that will serve as reference point in our novel simulations of the dynamics of this model quasicrystalline material. In particular, Sec. II introduces the LJG-type potential, outlines the phase diagram describing the dodecagonal quasicrystal system and provides other details regarding our computer simulations. In Sec. III, we quantify relaxation mechanisms, dynamic heterogeneity, bulk self-diffusion, and the intricate interplay among these basic dynamical properties. We observe temperature-dependent string-like motions within the quasicrystal. Finally, in Sec. IV, we conclude with a integrated discussion of our findings.
@@ II.1 Lennard-Jones-Gauss potential
We conducted classical MD simulations using the LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulation) software [39] to investigate the thermodynamic and dynamic properties of a 2D dodecagonal quasicrystal. Our system comprises identical particles interacting via a novel LJG-type potential, which combines the traditional Lennard-Jones potential with a secondary Gaussian well at longer distances. The LJG potential is defined as follows:
$V\left(r\right)=4{\varepsilon_{0}}\left[{{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-12}}-{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-6}}}\right]+\frac{h}{{{\sigma_{\rm{G}}}\sqrt{2\pi}}}\exp\left[{-\frac{{{{\left({r-{r_{\rm{G}}}}\right)}^{2}}}}{{2\sigma_{\rm{G}}^{2}}}}\right].$ (1)
In this equation, $\varepsilon_{0}$ is the energy minimum of the LJ well and $\sigma_{0}$ is defined as the zero-crossing distance for LJ potential. The parameter $h$, in conjunction with the standard deviation $\sigma_{\text{G}}$, dictates the peak height of the Gaussian component, while $r_{\mathrm{G}}$ denotes the peak position of the Gaussian well contribution to $V(r)$. Following previous work [30, 37, 34], the chosen parameter values that resulted in the formation of a dodecagonal quasicrystal are $\varepsilon_{\mathrm{0}}=0.5$, $\sigma_{\mathrm{0}}=0.88$, $h=-0.874$, $r_{\mathrm{G}}=1.879$, and $\sigma_{\textrm{G}}^{2}=0.02$. This choice is made by the prevalence of dodecagonal quasicrystals having this symmetry in real materials, especially soft matter materials such as polymers [33, 40, 41, 42]. All quantities are expressed in dimensionless Lennard-Jones units. We shall use temperature unit of $\varepsilon_{\mathrm{0}}/k_{\mathrm{B}}$ (here $k_{\mathrm{B}}=1$), energy unit of $\varepsilon_{\mathrm{0}}$, length unit of $\sigma_{\mathrm{0}}$ and time unit of $\tau=\sqrt{m\sigma_{\mathrm{0}}^{2}/\varepsilon_{\mathrm{0}}}$. A single MD time step is equal to $0.001\tau$. With these parameters, the LJG-type potential exhibits a double-well nature, as illustrated in Fig. 1(a). For practical purposes, we applied a cutoff to the LJG-type potential at $r=2.5$.
@@ II.2 Phase diagram
A system comprising $N=1024$ particles arranged in a square lattice is constructed. The system is heated to liquid phase in very fast rate and is fixed at $T=1.95$ finally. The simulations are conducted with periodic boundary conditions applied in both spatial dimensions. Following an initial period of 50,000 molecular dynamics (MD) steps to ensure equilibration, the system undergoes a controlled linear temperature decrease to reach $T=0.001$ over a course of $3.9\times 10^{7}$ MD steps. To maintain precise control over temperature and pressure within the isothermal-isobaric ensemble, all MD simulations are conducted in NPT ensemble which employ two crucial techniques. The Nosé-Hoover thermostat, as introduced by Nosé [43] and further refined by Hoover [44], regulates temperature fluctuations throughout the simulations. Meanwhile, the Parrinello-Rahman barostat, developed by Parrinello and Rahman [45], ensures stable pressure conditions. Throughout the paper the pressure is fixed at $P=0$. These simulations on the thermodynamic properties of our model are not meant to establish novel aspect of these materials, but rather to establish the thermodynamic and structural nature of the material as a reference pointy for understanding dynamical properties which are the novel aspect of the present work.
During the cooling process, the system exhibits two thermodynamic transitions. The transition from phase 1 (liquid) to phase 2 (hexatic) happens at about $T_{LH}=1.78$. In order to quantitatively identify the location of the transition between these phases, and qualify the nature of phase 2, we calculate the six-fold bond-orientation order parameter [46, 47] according to the relation,
@@
where $N_{j}$ is the number of nearest neighbors of particle $j$, $\mathbf{r}_{j}$ is the position of particle $j$, and $\theta_{m}^{j}$ is the angle between the vector ($\mathbf{r}_{m}-\mathbf{r}_{j}$) and the $x$ axis. The cutoff distance for determining these nearest neighbors is set to be $1.2$, which corresponds to the bond length derived from the radial distribution function (RDF), $g(r)$, as discussed below. The average order parameter is then obtained via $\overline{\Psi}_{6}=\frac{1}{N}\sum_{j=1}^{N}\psi_{6}^{j}$. The $T$ dependence of $\overline{\psi}_{6}$ is shown in Fig. 2(a), which clearly indicates two thermodynamic transitions at high and intermediate $T$.
@@
where $L$ is the length of the simulation box, and $N$ the number of atoms. The quasi-long-range bond-orientation order arises from the algebraic decay of $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-\eta_{6}}$. It is noteworthy that at the boundary between the liquid phase and the hexatic phase, $\eta_{6}$ assumes a value of 1/4. In the liquid phase, the system exhibits only short-range bond-orientational order. Consequently, the bond-orientation correlation function is more aptly described by the Ornstein-Zernike (OZ) function, characterized by $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-1/2}\exp\left(-r/\xi_{6}\right)$ [48, 49, 46]. As evident in Fig. 2(b), $g_{6}(r)/g(r)$ exhibits a more gradual decay with increasing temperature. We observe a power-law decay with an exponent of $\eta_{6}=-1/4$ for $g_{6}(r)/g(r)$ the characteristic temperature $T_{LH}=1.78$, providing a practical estimate of the location of the liquid-hexatic transition [48, 49, 46]. It is worth to mention a situation where quasicrystals melt without a hexatic phase. There exist situations in 2D, where quasicrystals transition into a square crystal, and then melt [50]. While quasicrystals are often high-temperature phases, it is hypothesized that there might exist quasi-crystalline ground states, the so-called ideal tilings [3, 4].
@@
Figure 3: The atomic structure of the monolayer dodecagonal (12-fold) quasicrystal, following relaxation at $T=0.3$. The dodecagonal unit and structural defect are highlighted by the circles.
The transition from phase 2 (hexatic) to phase 3 (quasicrystal) takes place near the characteristic temperature $T_{HQ}=0.87$. Through a detailed comparison of the local configurations in phase 3 with experimental observations obtained via high-resolution transmission electron microscopy (TEM) [51], as illustrated in Fig. 3, we confirm that phase 3 corresponds to a monolayer dodecagonal (12-fold) symmetry quasicrystal, pointing to the relevance of our simulations to real-world quasicrystal material. Our model quasicrystal exhibits dodecagonal units and a point defect structure within a hexagonal ring configuration, both of which align with the observations made through TEM [51] and other experimental observation [52]. Here, the number of defects can be controlled according to the cooling rate, but they are indeed unavoidable. This situation also exists in real quasicrystal materials [53]. The agreement between simulation results and experimental observations confirms the fidelity of our model in capturing the intricate atomic arrangement of the quasicrystal phase. The specific quasicrystal here is the so-called square-triangle tiling (STT) in the literature [54]. Actually, it is atypical for quasicrystals, and even atypical for dodecagonal quasicrystals, due to the fact that it does not have phason modes. Instead, its dynamics is dominated by the diffusion of semi-defects (or partial dislocations) as observed in Fig. 3, a process called Zippers motion [55].